the distance of an object traveling is a straight line is s=t^2-8t where s is in feet and t is in seconds. how long will take for an object to travel 9ft?
Answer by mananth(16946)

the distance of an object traveling is a straight line is s=t^2-8t where s is in feet and t is in seconds. how long will take for an object to travel 9ft?
s=9
9=t^2-8t
t^2-8t-9=0
t^2-9t+t-9=0
t(t-9)+1(t-9)=0
(t-9)(t+1)=0
t=9 or -1
After 9 seconds since t=-1 is not possible